MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2007/07/12
OBSERVER: Fukui Furusawa Pam
CONDITION: clear
COMMENT: Dark D2451(60sec), D2452(300sec) are OK. An anomaly seemed to be found in MOA305 on gb1 and we observed this intensively, but it was wrong idea.
Compressor gas pressure: 1.8 MPa
Expected Observation Time: START 18:30 END 07:16


TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB8591-871612638723220
LMC1590-15901123600
SMC396-39612600
ALERT-000
GRB-000
TOTAL---12840127420 (07h37m00s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
831312115110111052101411
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
1010101010101099110


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
111101111011
Field13Field14Field15Field16
1100


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
110000000000


SPECIAL TARGET NAMEBR>
CategoryTargetNrun


CHECK SHEET

TimeStatusWeather
17:00waitingclear. Mirror was frosted, lenses were fogged.
18:35observingclear
19:00observingclear
20:00observingclear
21:00observingclear
22:00observingclear, we observed only gb1 for MOA305
23:00observingclear, we observed only gb1 for MOA305
00:00observingclear
01:00observingclear
02:00observingclear
03:00observingclear
04:00observingclear
05:00observingclear
06:00observingclear
07:10observingclear, finish observation


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day